3. Calculate the mass (in grams) of calcium nitrate [Ca(NO3)2] that can be produced by reacting 136 mL of 4.0 M nitric acid (HNO3)2 with excess calcium hydroxide [Ca(OH)2].
(molar mass of Ca(NO3)2 = 164.1g/mol)
1
Expert's answer
2017-11-04T14:56:07-0400
Ca(OH)2 + 2HNO3 = Ca(NO3)2 + 2H2O
n(HNO3) = c × V = 0.136 × 4 = 0.544 mol
n(HNO3):n(Ca(NO3)2) = 2:1
n(Ca(NO3)2) = n(HNO3)/2 = 0.544/2 = 0.272 mol
m(Ca(NO3)2) = n(Ca(NO3)2) × M(Ca(NO3)2) = 0.272 × 164.1 = 44.64 g